The maximum value of sin(x + p / 6) + cos(x + p / 6) in the interval (0, p / 2) is attained at x =

p / 2
p / 3
p / 6
p / 12

Grade:12

1 Answers

bharat bajaj IIT Delhi
10 years ago
sin(x + p / 6) + cos(x + p / 6) =Tignometry-Sum and Difference formulae
=Trigmometric ratios
=Trigonometric ratio

This is maximumwhen sin(x + 5p / 12) ismaximum.
That is when sin(x + 5p / 12) = 1
x + 5p / 12 = p / 2
x = p / 12
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty